OpenStudy (anonymous):

In the class elections, Connie received 53 more votes than Carl. In all, 221 votes were cast for the two candidates. How many votes did each receive

OpenStudy (jeyekomon):

And what exactly is difficult to solve on it?

OpenStudy (anonymous):

Carl gor 84 and Connie got 137

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

how would I set the equation up

OpenStudy (anonymous):

Let carl got \( x\) votes, then Connie got \( 53 + x \) votes. SO, \( 2x + 53 = 221 \Rightarrow x= \frac{168}{2} =84\) and hence, connie got 137.

OpenStudy (anonymous):

thank you again

OpenStudy (anonymous):

let carl be x then connie will x+53 x+x+53=221 2x=168 x=84

OpenStudy (jeyekomon):

Wow, I admire you, FoolForMath, for using the equation tool. It would take ages for me to write that, so I rather use just plaintext..